OpenStudy (anonymous):

A train leaves the station at 9:00 pm traveling at 36 miles per hour. A second train leaves the station at 10:00pm traveling at 42 miles per hour. At what time will both trains have traveled the same distance? A.) 4:00 am B.) 5:00 am C:) 6:00 am D:) 7:00

OpenStudy (baby_bear69):

I could figure it up and give you the answer, but I do it the difficult way.. Idk how to explain

OpenStudy (anonymous):

I really need the answer I've been stuck for awhile. :/ please help!

OpenStudy (baby_bear69):

Basically im just making a chart. and as you add an hour, add 36 or 42 cause thats how many miles they go per hour

OpenStudy (anonymous):

Okay thanks :/ I tried that but I think I messed up

OpenStudy (baby_bear69):

okok i got it

OpenStudy (baby_bear69):

Its A

OpenStudy (baby_bear69):

can i get a medal please, just hit Best Response

OpenStudy (radar):

Just in case you are interested in a method to solve this: Distance = Rate (mph) times time (hours) or D=RT This is the relationship you use. For this problem, the distance traveled by first train will equal the distance traveled by second train.

OpenStudy (radar):

D (distance) of first train = speed of first train times the time that the train traveled D of 2nd train = speed of 2nd train times the time traveled by 2nd train. The unknown here is the times. We now resort to algebra.

OpenStudy (anonymous):

picking up from where @radar left off 

OpenStudy (anonymous):

36t = 42(t - 1) 36t = 42t - 42 -6t = -42 t = 7 hours the trains have traveled the same distance 7 hours after the first train leaves the station. That's 4:00 am

OpenStudy (radar):

Let X = the amount of time traveled by first train. Since 2nd train left an hour later .....then..... X-1 = the amount of time traveled by 2nd train. Since the distance traveled at the point where they were at the same place is the same. 36X = 42(X-1) Now all that is required is to solve for x. Add the value obtained for x to the time that first train left and you will have the time where both trains traveled the same distance. PS For some reason there was a connection failure.

OpenStudy (radar):

Let X = the amount of time traveled by first train. Since 2nd train left an hour later .....then..... X-1 = the amount of time traveled by 2nd train. Since the distance traveled at the point where they were at the same place is the same. 36X = 42(X-1) Now all that is required is to solve for x. Add the value obtained for x to the time that first train left and you will have the time where both trains traveled the same distance. PS For some reason there was a connection failure.

OpenStudy (radar):

However, I see that @peachpi has provided an excellent solution.